A 11.8-kg cylinder rolls without slipping on a rough surface. At an instant when its center of gravity has a speed of 11.1 m/s, determine the following.

(a) the translational kinetic energy of its center of gravity
J

(b) the rotational kinetic energy about its center of gravity
J

(c) its total kinetic energy
J
Answer #1

(a) the translational kinetic energy of its center of mass


(1) KEt =
Answer #2

a)1/2 mv^2=0.5*11.8*11.1^2=726.939 j
b)I = (1/2) m r^2
w = (v/r)
(for a rolling object, the velocity of the center of mass is equal to the tangential velocity, I think)

KErot = (1/2) (1/2) m r^2 (v/r)^2
KErot = (1/4) m v^2
(That is one half of the translational KE or 1/2 the answer of part a)


=363.469 joule


c) a+b=1090.40 joule